OpenStudy (anonymous):

-5i(-3+5i) tell me how you got your answer

OpenStudy (anonymous):

use commutative law -5i(-3) + (-5i)(5i) = 15i + 25

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

15i+25

OpenStudy (anonymous):

Distributive property. -5i * -3 = 15i -5i*5i= -25i^2 -25i^2 + 15i

OpenStudy (anonymous):

lovelesstime i^2 cancels out

OpenStudy (anonymous):

i^2=-1

OpenStudy (anonymous):

Two i's. it's i to the second power...

OpenStudy (anonymous):

Dheera am I right?

OpenStudy (anonymous):

yes i know \[i^2=-1\]

OpenStudy (anonymous):

It's 25i^ + 15i...

OpenStudy (anonymous):

\[15i+25\] is the correct answer. i was just double checking. Thanks.

OpenStudy (anonymous):

use i^2=-1

OpenStudy (anonymous):

No..it's not that. It's 5i * 5i There is 2i's so 25i^2

OpenStudy (anonymous):

YOU ARE RIGHT BUT \[I=\sqrt{-1} \] SO I^2=-1

OpenStudy (anonymous):

no its lower case i = sqrt(-1)

OpenStudy (anonymous):

I didn't know that. -.- I look at it and see 25i^2. I don't know that i ever equaled. I'm 14.

OpenStudy (anonymous):

wts the confusion about?

OpenStudy (anonymous):

\[\sqrt{-1}=i \sqrt{1}\]

OpenStudy (anonymous):

It's 25i^2 + 15i

OpenStudy (anonymous):

no. it's \[25+15\]

OpenStudy (anonymous):

25+15i

OpenStudy (anonymous):

You forgot the i!

OpenStudy (anonymous):

i fixed it.

OpenStudy (anonymous):

Right now i is a variable! it is 25i^2 + 15i

OpenStudy (anonymous):

in the question it does not give i!

OpenStudy (anonymous):

i dont know what you're asking for me for Em, you are smashing this question to obvilion! it doesn't have a chance agains you!

OpenStudy (anonymous):

sorry Lovelesstime got to go. no time to explain. but since i=imaginary number. http://en.wikipedia.org/wiki/Imaginary_number little more explaining in wiki.

OpenStudy (anonymous):

It is 25i^2 5i * 5i = 25i^2 they do not give i in the question therefore I am right. i is a variable in this question not -1 because in the question it does not SAY it is.

OpenStudy (anonymous):

mostly i should taken as imaginary number.

OpenStudy (anonymous):

are you doing complex numbers or not? if you are then i is the square root of -1

OpenStudy (anonymous):

i could be anything!

OpenStudy (anonymous):

I've never seen i used as a variable

OpenStudy (anonymous):

I have.

OpenStudy (anonymous):

i mean while doing complex numbers

OpenStudy (anonymous):

i is the imaginary unit, UNLESS SPECIFIED...finished

OpenStudy (anonymous):

This means I'm right, though.

OpenStudy (anonymous):

ur right in ur own way emily, but its just that u didnt know what i was

OpenStudy (anonymous):

The question did not say.

OpenStudy (anonymous):

no question says it

OpenStudy (anonymous):

sexy

OpenStudy (anonymous):

It didn't say complexed o2 anythin.

OpenStudy (anonymous):

or*

OpenStudy (anonymous):

keep chugging away guys, ill help soon as i finish on anoterh quesiotn

OpenStudy (anonymous):

arre yaar.....y dont u understand...,i is used only for complex numbers,,they dont use it anywhere else

OpenStudy (anonymous):

rofl@akbar

OpenStudy (anonymous):

Yes, they do. I see them in my math textbook all the time.

OpenStudy (anonymous):

I am not clicking.

OpenStudy (anonymous):

its safe for work, dont worry

OpenStudy (anonymous):

thats juvenile

OpenStudy (anonymous):

somebody calm her down..she isnt ready to accept the standard notation for iota

OpenStudy (anonymous):

I am right gimme my medal Paul. -.-

OpenStudy (anonymous):

meh I'm 17, I can be juvenile if I want

OpenStudy (anonymous):

are you 17 loveless time? im 25 :( so old, and so much math to learn!

OpenStudy (anonymous):

I'm 14.

OpenStudy (anonymous):

who gave medals for lovelesstime. did they give to calm her down